OpenStudy (baldymcgee6):

Difference between negative exponent and inverse functions? For example, sine inverse: sin^-1(x) or 1/(sin(x)) = (sin(x))^-1

OpenStudy (campbell_st):

the negative indice means to find the reciprocal of sin(x) so given \[\sin(x) = \frac{\sin(x)}{1}\] then the reciprocal is \[(\frac{\sin(x)}{1})^{-1} = \frac{1}{\sin(x)}\] and for the inverse function if \[\sin(x) = \frac{a}{b}\] the sin of an angle is equal to a ratio then \[x = \sin^{-1}(\frac{a}{b})\] the angle is equal to the inverse sin of the ratio. its the 2 way connection between and angle and a ratio. hope it helps.

OpenStudy (baldymcgee6):